OpenStudy (anonymous):

Graph the equation and identify the y intercept y=1/4 x

OpenStudy (amistre64):

its a flat line that hits the origin

OpenStudy (amistre64):

well...... flattish

OpenStudy (anonymous):

is it \[\frac{1}{4x}\] or \[\frac{1}{4}x\]

OpenStudy (anonymous):

The y intercept is at x = 0, as this is where the line crosses the y-axis.

OpenStudy (anonymous):

the first one does not cross the y - axis. the second one crosses at (0,0) it is a line with slope \[\frac{1}{4}\] through the origin

OpenStudy (anonymous):

1/4 x

OpenStudy (anonymous):

amistre, like my new icon? it is a cute girl, done by mondrain

OpenStudy (anonymous):

she was that way when i found her officer...

OpenStudy (amistre64):

post deleted for posterities sake ;)

OpenStudy (anonymous):

r u going to help me or just chat on my page

OpenStudy (amistre64):

both actually; since the question was answered and you provided no feedback.... we chat

OpenStudy (anonymous):

how do I graph it on a graph puzzled on how

OpenStudy (amistre64):

think of it as squatting in the rom inthe hopes that you will return :)

OpenStudy (amistre64):

you start at zero and plot another point at let say 4,1 and another at 8,2 and conect the dots

OpenStudy (anonymous):

thank u.. that is what I was looking for.

OpenStudy (amistre64):

the basic step is ust make a table of points ; whatever you use for x, makes a y

OpenStudy (amistre64):

:) youre welcome

OpenStudy (anonymous):

sorry it has been a long time since i have taken a math class it is a little fuzzy

OpenStudy (amistre64):

itll come back to ya ;)

OpenStudy (anonymous):

It has , but the graphing is confusing me I am not understanding it

OpenStudy (amistre64):

x | y = (1/4)x ----------- -1 | -1/4 0 | 0 1 | 1/4 2 | 1/2 4 | 1